OpenStudy (anonymous):

Solve. 3x - 5 = 7x - 17

OpenStudy (anonymous):

add 17 to both sides. 3x +12 = 7x subtract 3x from both sides 12 = 4x divide by 4 and you'll have what x equals

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

3x-5+5=7x-17+5 3x=7x-12 3x-7x=7x-12-7x -4x=-12 divide both sides by -4 and you will get x=3